A simple pendulum has a period of 4.2 secs. When the length of the pendulum is shortened by 1 meter, the period is 3.7 sec. Calculate the original length. The value of acceleration due to gravity.
1
Expert's answer
2019-03-22T10:34:00-0400
"T_1=2\pi \sqrt{\frac{l}{g}}"

"T_2=2\pi \sqrt{\frac{l-x}{g}}"

"\frac{l-x}{l}=(\frac{T_2}{T_1})^2"

"\frac{l-1}{l}=(\frac{3.7}{4.2})^2""l = 4.5 m.""g=l(\frac{2 \pi}{T_1})^2=4.5(\frac{2 \pi}{4.2})^2=10 m/s^2"